prove that
2sin^2theta+2cos^2theta-1=0
1
Expert's answer
2020-09-20T18:06:24-0400

"2\sin^2\theta + 2\cos^2\theta -1 = 2*(sin^2\theta + \cos^2\theta) - 1 = 2-1 =1"

for any value "\theta"

hence the assumption "2\sin^2\theta + 2\cos^2\theta -1 = 0" is not true


Answer: The original hypothesis requiring proof is false